OpenStudy (anonymous):

x=sqrt(7+4sqrt3) then (x+(1/x))=?

OpenStudy (anonymous):

\[ x + \dfrac {1}{x} = \dfrac {x^2 + 1}{x} \]You can use that. Or: You could just do it the regular way. You would do \[ \dfrac {1}{x} = \dfrac {1}{7+4\sqrt{3}} = \frac {7-4\sqrt{3}}{(7+4\sqrt{3})(7-4\sqrt{3})}. \]

OpenStudy (anonymous):

sry frnd but it's become too deficult

OpenStudy (anonymous):

so it is wrong ans.

OpenStudy (raden):

4

OpenStudy (anonymous):

right ans.

OpenStudy (raden):

do u know how come it ?

OpenStudy (anonymous):

yes

OpenStudy (raden):

hehe,, good

OpenStudy (anonymous):

but you looking smart person

OpenStudy (raden):

sometimes ^^

OpenStudy (anonymous):

Dude, you have to read everything someone writes. I didn't finish the solution. I only gave a hint.